OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that has the given slope and contains the given point: m= 2/3 (9,5) Write an equation in slope-intercept form for the line that has the given slope and contains the given point: m= 2/3 (9,5) @Mathematics

OpenStudy (anonymous):

y = mx + b First, put in the slope y = 2/3 x + b Next put in the x and y values and solve for b 5 = (2/3) 9 + b 5 = 6 + b b = -1 y = 2/3 x -1